OpenStudy (anonymous):

How would you graph this equation: x-4y=8

OpenStudy (anonymous):

would u just minus the x to the other side and end up with -4y=-x+8?

OpenStudy (anonymous):

then you have to divide by -4. do you know what to do from there?

OpenStudy (anonymous):

oh, hold on

OpenStudy (anonymous):

isolate the y. x= 8+4y, then x-8=4y, then (x-8)/4=y, and sometimes it is easier to see in standard y=mx+b so y=x/4-2=y. I'll let you do the algebra to figure how i simlified it.

OpenStudy (anonymous):

now you can plot y vs x

OpenStudy (anonymous):

after you divide the -4y to both sides dont you end up with: y= -1/4x+2

OpenStudy (anonymous):

*-2

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

you could do this. the way you started would give -4y=8-x then y=(8-x)/(-4 ), then y = -2 +x/4 = y. same result :). good job.

OpenStudy (anonymous):

thanks guys! :D

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

im going to draw the graph. hold on

OpenStudy (anonymous):

b is the intercept and m is the slope. in this case we have a negative intercept and a positive slope. good luck. gtg.